OpenStudy (anonymous):

Which equation matches the sentence? Column A Column B 1. Seven more than 2 times a number is 23. 2. The difference of 2 times a number and 7 is 23. 3. The sum of 7 times a number and 2 is 23. 4. Seven less than 2 times a number is 23.

Mehek (mehek14):

where are the equations?

OpenStudy (anonymous):

column b is a. 7n + 2=23 b. 2n - 7=23 c. 2n + 7=23

OpenStudy (anonymous):

these are the equations

OpenStudy (anonymous):

a. goes with 3.

OpenStudy (anonymous):

okay what esle

OpenStudy (anonymous):

b. with 4.

OpenStudy (anonymous):

and

OpenStudy (anonymous):

and 1 goes with what

OpenStudy (anonymous):

c

OpenStudy (anonymous):

so 2 with

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanxz